# Non-negative polynomials on generalized elliptic curves

### v2

@@ 1. Introduction
While in the smooth case 1.1 is a direct consequence of the Riemann–Roch theorem, as observed for very ample $L$ in [baldi], the singular case is considerably more difficult. A consequence of 1.1 is that every non-negative global section on a generalized elliptic curve is a sum of squares of global sections of a not necessarily invertible sheaf:
@@
A direct consequence of 1.2 is that the dual cone $\operatorname{P}(X,L)^{\vee}$ can be described as the cone of all linear forms $\ell\colon H^{0}(X,L\otimes L)\to{\mathbb{R}}$ such that
@@
Spectrahedra are convex semi-algebraic sets defined by a linear matrix inequality. They are fundamental objects in convex algebraic geometry; they are the feasible sets of semi-definite programming and have many desirable properties [clausbook, timdaniel, siambook]. Note that, while the convex hull of a curve is always a spectrahedral shadow [scheiderercurves, scheiderercurves2], i.e., the image of a spectrahedron under a linear map, it is rarely a spectrahedron itself. A more careful analysis of the proof of 1.3 (see 3.8) will show that the description of the convex hull of $X({\mathbb{R}})$ in 1.3 as a spectrahedron is particularly nice in that it is given by block matrices with blocks of size at most $\frac{n+1}{2}$. Convex hulls of elliptic curves were studied before in [scheiderergenusone, kummersinn].
In order to deduce 1.3 from 1.2, we need a certain divisibility result on the Picard group of a generalized elliptic curve. In Section 2.3 we prove such for arbitrary real curves. From this we also deduce a generalization to singular curves of a result by Geyer–Martens [geyermartens] on $2$-torsion points in the Picard group of a smooth real curve, which might be of independent interest (2.3).
@@ 2.3. The Picard group of real curves
As the support of $D$ does not contain real points, we can write its pullback to $X_{\mathbb{C}}$ as $F+\bar{F}$ for some divisor $F$ where $\bar{F}$ denotes its complex conjugate. Let $F_{1}$ be a divisor supported on regular points fixed by the complex conjugation such that ${\mathcal{O}}_{X_{\mathbb{C}}}(F-F_{1})\in\operatorname{Pic}^{\circ}(X_{\mathbb{C}})$. Since $\operatorname{Pic}^{\circ}(X_{\mathbb{C}})$ is $2$-divisible by 2.1, there exists a divisor $G$ supported on regular points of $X_{\mathbb{C}}$ and a unit $g$ of the total quotient ring of $X_{\mathbb{C}}$ such that $2G+\operatorname{div}(g)=F-F_{1}$. This implies that
@@
2.2 can be used to generalize a result by Geyer–Martens [geyermartens] on the structure of $2$-torsion points on a smooth real curve to singular curves. Namely, for a connected curve $X$ over $\mathbb{R}$} we consider the map
@@
Because $X$ is projective, there exists an open affine subset $U$ of $X$ which contains $X({\mathbb{R}})$ and all singular points. Indeed, if $X$ is embedded to $\mathbb{P}^{n}$, then, after applying a linear change of coordinates if necessary, we can take the principal open subset of all points where $x_{0}^{2}+\cdots+x_{n}^{2}$ is non-zero. Embedding $U$ to ${\mathbb{A}}^{m}$ gives an embedding of $X({\mathbb{R}})$ to ${\mathbb{R}}^{m}={\mathbb{A}}^{m}({\mathbb{R}})$. Let $X_{1}$ be a union of connected components of $X({\mathbb{R}})$ and $X_{2}=X({\mathbb{R}})\smallsetminus X_{1}$. Since $X_{1}$ and $X_{2}$ are disjoint compact subsets of ${\mathbb{R}}^{m}$, there is a continuous function $\psi\colon{\mathbb{R}}^{m}\to{\mathbb{R}}$ which satisfies $\psi(x)\leq-1$ for $x\in X_{1}$ and $\psi(x)\geq 1$ for $x\in X_{2}$. By the Stone–Weierstrass theorem we can uniformly approximate $\psi$ on a compact subset of ${\mathbb{R}}^{m}$ containing $X({\mathbb{R}})$ arbitrarily close by a polynomial. In particular, there exists $g\in{\mathbb{R}}[x_{1},\ldots,x_{m}]$ with $g(x)<0$ for $x\in X_{1}$ and $g(x)>0$ for $x\in X_{2}$. After replacing $g$ by a small perturbation if necessary, we can further assume that $g$ does not vanish on any singular point of $X$. We apply 2.2 to the principal divisor $D=\operatorname{div}(g)$ and obtain a divisor $E$, supported on regular points, and a unit $f$ of the total quotient ring of $X$, which is non-negative on every real point where it is defined, such that $\operatorname{div}(\frac{g}{f})=2E$. The $2$-torsion point ${\mathcal{O}}_{X}(E)$ of $\operatorname{Pic}(X)$ is then mapped by $\rm sg$ to a cohomology class represented by the function which is 1-1 on $X_{1}$ and +1+1 on $X_{2}$. Since $X_{1}$ was chosen to be an arbitrary union of connected components of $X({\mathbb{R}})$ and $X_{2}$ its complement, this shows the claim. ∎
We will see that positive $2$-torsion points are closely related to those extreme rays of $\operatorname{P}(X,L)$ with only regular zeros. 2.3 gives a way to count them.
@@ 2.4. Generalized elliptic curves
Here the first equality is a classical fact about real elliptic curves and the second equality holds by 2.3. This proves that $\dim_{{\mathbb{F}}_{2}}(\operatorname{Pic}(X)^{+}_{2})=1$. ∎
@@ 3. Proofs and examples
The dimension of the image of $H^{0}(\tilde{X}_{i},M_{i})$ under $\beta$ is bounded from above by the sum $\delta_{i}=\sum_{x\in X_{i}}\delta(x)$. Using 2.7 and going through 2.6 we see that this sum is at most $1$ for irreducible $X$ and that in general $\delta_{i}\leq 2$ for every $i$. This property can also be deduced directly from the definitions by making use of the fact that $\pi^{!}_{i}\omega^{\circ}_{X}$, where $\pi_{i}\colon\tilde{X}_{i}\to X$ is the inclusion $\tilde{X}_{i}\hookrightarrow\tilde{X}$ composed with the normalization $\pi\colon\tilde{X}\to X$, is, on the one hand, the canonical sheaf of $\tilde{X}_{i}$ by [hartshorne]*Exercise III.7.2, and, on the other hand, the codifferent sheaf $\pi^{!}_{i}{\mathcal{O}}_{X}$ of $\pi_{i}$.
@@
If $X$ is smooth, then the claim follows from the Riemann–Roch theorem. Hence we assume from now on that $X$ is singular. Then each $\tilde{X}_{i}$ is isomorphic to $\mathbb{P}^{1}$ and thus the space $W$ of all $h\in H^{0}(\tilde{X}_{j},M_{j})$ with $\operatorname{div}(h)\geq\operatorname{div}(f_{j})-p-q$ has dimension $3$. By 3.1 the dimension of $\beta(H^{0}(\tilde{X}_{j},M_{j}))$ is at most $2$. Since $\dim(W)=3$, there is $0\neq h\in W\cap\ker(\beta)$. Thus the claim follows if $h$ is linearly independent from $\alpha(f)$. Hence we assume in the following that $h$ and $\alpha(f)$ are linearly dependent. This can only happen if $f_{i}=0$ for all $i\neq j$ and thus $\beta(f_{j})=0$. If $r>1$, then for $g\in H^{0}(\tilde{X}_{j},M_{j})$ the condition $\beta(g)=0$ is equivalent to a vanishing condition on the points on $\tilde{X}_{j}$ that are mapped by $\pi$ to some $X_{i}$ with $i\neq j$. Since $f_{j}$ satisfies this condition and since neither $p$ nor $q$ are mapped by $\pi$ to some $X_{i}$ with $i\neq j$, the condition $\operatorname{div}(g)\geq\operatorname{div}(f_{j})-p-q$ implies that $g\in H^{0}(\tilde{X}_{j},M_{j})$ satisfies this vanishing condition as well. Thus $W$ lies in the kernel of $\beta$ which implies the claim. It remains to treat the case that $r=1$. Then by 3.1 the dimension of $\beta(H^{0}(\tilde{X}_{j},M_{j}))$ is $1$ and thus the space $W\cap\ker(\beta)$ is at least of $2$-dimensional which implies the claim. ∎
Proof of 1.1.
@@
Let $X_{j}$ be an irreducible component of $X$ that contains $x$. Note that this implies $j\not\in S$. There exists $p\in\pi^{-1}(x)\cap\tilde{X}_{j}$ and we let $q=\bar{p}$ be its complex conjugate. Because $x\not\in X({\mathbb{R}})$, the points $p$ and $q$ are two distinct elements of $\tilde{X}_{j}$. Now $f$, $p$ and $q$ satisfy the assumptions of 3.2. Thus there exists $g\in H^{0}(X,L\otimes L)$ which is linearly independent of $f$ such that:
@@
Proof of 1.2.
Let $D$ be a divisor on $X$ supported on regular points such that $L\cong{\mathcal{O}}_{X}(D)$, the subsheaf of the sheaf of total quotient rings of ${\mathcal{O}}_{X}$ associated to $D$. Let $s\in H^{0}(X,L\otimes L)$ generate an extreme ray of $\operatorname{P}(X,L)$. Consider the partial normalization $\tilde{Y}\to X$ at all singular points of $X$ where $s$ vanishes. Further, let $Y\subseteq\tilde{Y}$ be the union of all irreducible components of $\tilde{Y}$ where the pullback of $s$ to $\tilde{Y}$ is not identically zero. Finally, let $\rho\colon Y\to X$ be the composition of the inclusion $Y\hookrightarrow\tilde{Y}$ with $\tilde{Y}\to X$. Then the restriction $D^{\prime}=D|_{Y}$ satisfies $\rho^{*}L\cong{\mathcal{O}}_{Y}(D^{\prime})$. We choose $f\in{\mathcal{O}}_{Y}(2D^{\prime})$, a rational function on $Y$, which corresponds to $\rho^{*}s$ under $\rho^{*}L\cong{\mathcal{O}}_{Y}(D^{\prime})$. We denote by $E_{1}$ the restriction of the zero divisor of $\rho^{*}s$ to the set of all regular points of $Y$ that lie over a singular point of $X$. Then the principal divisor of $f$ is given by $\operatorname{div}(f)=E_{1}+E_{2}-2D^{\prime}$ for some effective divisor $E_{2}$ supported on regular points. By 1.1 the divisor $E_{2}$ is a sum of real non-isolated points of $Y$. Because $f$ is non-negative on $Y({\mathbb{R}})$, it follows that $E_{2}=2F$ for some divisor $F$ because real zeros of non-negative sections must occur with even multiplicity. Consider the bilinear map
@@
The proof of 1.2 simplifies when the extreme ray $s$ has only regular zeros. In this case, the proof shows that $s$ is the image of a square under a positive semi-definite isomorphism $\psi\colon A\otimes A\to L\otimes L$ where $A$ is a line bundle. Hence we have $A=L\otimes T$ where $T\in\operatorname{Pic}(X)_{2}$ is an positive $2$-torsion point. By 2.8 there are at most two such points. However, as 3.5 shows, line bundles are not enough to describe extreme rays that vanish at singular points.
@@
Consider the case that $X$ is a smooth elliptic curve over $\mathbb{R}$}. A plane affine model of $X$ is given by the equation $y^{2}=x\cdot p(x)$ where $p$ is a univariate polynomial of degree $2$ with simple zeros which are either positive or non-real. Let $P$ be the point $(0,0)$ and let $O$ be the point at infinity. Then the unique (see 2.8) non-trivial positive $2$-torsion point is represented by the divisor T=POT=P-O. Indeed, we have $\operatorname{div}(x)=2T$ and $x$ is non-negative on $X({\mathbb{R}})$. Hence, if $D$ a divisor on $X$ and $L={\mathcal{O}}_{X}(D)$, then every extreme ray of $\operatorname{P}(X,L)$ is either a square of a global section of ${\mathcal{O}}_{X}(D)$ or the image of a square under the map
@@
We illustrate the proof of 1.2 in the case when $X$ is the rational curve with one cusp. A plane affine model of $X$ is given by the equation $y^{2}=x^{3}$. Let $O$ be the point at infinity and $L={\mathcal{O}}_{X}(dO)$ for $d\in{\mathbb{N}}$. Note that for $d=1,2$ the line bundle $L$ is ample but not very ample. The only possibilities for $\rho\colon Y\to X$ are the identity $X\to X$ and the normalization $\mathbb{P}^{1}\to X$. In the first case, we have $E_{1}=0$. Since the Picard group of $X$ is the additive group $\mathbb{R}$$\times{\mathbb{Z}}$, which is torsion-free, the only possibility, up to a positive scalar, for $\psi$ is the identity on $L\otimes L$. In the case $Y=\mathbb{P}^{1}$, we denote by $P$ the preimage of the cusp under $\pi$. Since $E_{2}$ has even degree and $E_{1}$ is linearly equivalent to $2dO-E_{2}$, we either have $E_{1}=2kP$ for $k=1,\ldots,d$. Here, by abuse of notation, we denote the preimage of $O$ under $\pi$ also by $O$. Again, as the Picard group of $\mathbb{P}^{1}$ is torsion-free, we find that $F$ is linearly equivalent to $dO-kP\sim(d-k)O$. In order to describe the resulting map $\psi$ on global sections explicitly, we identify the global sections of ${\mathcal{O}}_{X}(mO)$ and ${\mathcal{O}}_{Y}(mO)$ with the spaces ${\mathbb{R}}[t^{2},t^{3}]_{\leq m}$ and ${\mathbb{R}}[t]_{\leq m}$ of univariate polynomials in ${\mathbb{R}}[t^{2},t^{3}]$ and ${\mathbb{R}}[t]$ that have degree at most $m$. We get the map
@@
1.2 gives the following nice description of the dual cone $\operatorname{P}(X,L)^{\vee}$.
@@
The proof of 1.2 shows that the description of $\operatorname{P}(X,L)^{\vee}$ in 3.6 consists of several blocks, one corresponding to each of the sheaves $A_{1},\ldots,A_{m}$ on $X$. A more careful analysis gives bounds on the size of these blocks. By the construction, the size of the block corresponding to $A_{i}$ is equal to $h^{0}(X,A_{i})$. Using the notation as in the proof of 1.2, this number is equal to $h^{0}(Y,{\mathcal{O}}_{Y}(F))$. By the Riemann–Roch theorem [stacks-project]*Lemma 0BS6 we have
@@
The proof of 1.3 is a combination of 3.6 and $2.$
Proof of 1.3.
Because $X({\mathbb{R}})\subseteq{\mathbb{R}}^{n}=(\mathbb{P}^{n}\smallsetminus H)({\mathbb{R}})$ is compact, there exists an affine linear functional ${\mathbb{R}}^{n}\to{\mathbb{R}}$ which is positive on $X({\mathbb{R}})$. Let $\ell$ be the corresponding global section of $M={\mathcal{O}}_{X}(1)$. By slightly perturbing $\ell$ if necessary, we can assume that the zero set of $\ell$ consists only of regular points of $X$. Let $D$ be its zero divisor. By 2.2 there exists a divisor $E$, supported on regular points, and a unit $f$ of the total quotient ring of $X$, which is non-negative on every real point where it is defined, such that $D=2E+\operatorname{div}(f)$. We have isomorphisms
@@
that maps a point to the corresponding point evaluation, satisfies ${\rm ev}^{-1}(K^{\vee})=\operatorname{conv}(X({\mathbb{R}}))\subseteq{\mathbb{R}}^{n}$. Therefore, the set $\operatorname{conv}(X({\mathbb{R}}))={\rm ev}^{-1}(\psi^{\vee}(\operatorname{P}(X,L)^{\vee}))$ is a spectrahedron by 3.6. ∎
@@
It follows from 3.7 that, in the situation of 1.3, the description of $\operatorname{conv}(X({\mathbb{R}}))\subseteq{\mathbb{R}}^{n}$ consists of blocks, each of size at most $\frac{n+1}{2}$. Note that in general one cannot expect a description that only involves blocks of smaller size, even when allowing a description as a spectrahedral shadow [scheiderercurves2].
@@
One checks that $X$ is a smooth elliptic curve. The hyperplane defined by $x_{0}=0$ at infinity does not intersect $X({\mathbb{R}})$. Thus 1.3 says that the convex hull of $X({\mathbb{R}})$ in the affine chart $x_{0}\neq 0$ is a spectrahedron. Indeed, it is the set of all $(x_{1},x_{2},x_{3})$ such that the matrix
@@
Therefore, it is isomorphic to the curve from 3.5 and ${\mathcal{O}}_{X}(1)$ is isomorphic to ${\mathcal{O}}_{X}(4O)$ where $O=(1:0:0:0)$. We use the notation as in 3.5. The hyperplane defined by $x_{0}=0$ at infinity does not intersect $X({\mathbb{R}})$. Thus 1.3 says that the convex hull of $X({\mathbb{R}})$ in the affine chart $x_{0}\neq 0$ is a spectrahedron. Indeed, it is the set of all $(x_{1},x_{2},x_{3})$ such that the matrix
@@
Figure 1. The (generalized) elliptic curves from 3.9 (left) and 3.10 (right).

### v3

@@ 1. Introduction
While in the smooth case Theorem 1.1 is a direct consequence of the Riemann–Roch theorem, as observed for very ample $L$ in [baldi], the singular case is considerably more difficult. A consequence of Theorem 1.1 is that every non-negative global section on a generalized elliptic curve is a sum of squares of global sections of a not necessarily invertible sheaf:
@@
A direct consequence of Theorem 1.2 is that the dual cone $\operatorname{P}(X,L)^{\vee}$ can be described as the cone of all linear forms $\ell\colon H^{0}(X,L\otimes L)\to{\mathbb{R}}$ such that
@@
Spectrahedra are convex semi-algebraic sets defined by a linear matrix inequality. They are fundamental objects in convex algebraic geometry; they are the feasible sets of semi-definite programming and have many desirable properties [clausbook, timdaniel, siambook]. Note that, while the convex hull of a curve is always a spectrahedral shadow [scheiderercurves, scheiderercurves2], i.e., the image of a spectrahedron under a linear map, it is rarely a spectrahedron itself. A more careful analysis of the proof of Corollar 1.3 (see Remark 3.8) will show that the description of the convex hull of $X({\mathbb{R}})$ in Corollary 1.3 as a spectrahedron is particularly nice in that it is given by block matrices with blocks of size at most $\frac{n+1}{2}$. Convex hulls of elliptic curves were studied before in [scheiderergenusone, kummersinn].
In order to deduce Corollary 1.3 from Theorem 1.2, we need a certain divisibility result on the Picard group of a generalized elliptic curve. In Section 2.3 we prove such for arbitrary real curves. From this we also deduce a generalization to singular curves of a result by Geyer–Martens [geyermartens] on $2$-torsion points in the Picard group of a smooth real curve, which might be of independent interest (Corollar 2.3).
@@ 2.3. The Picard group of real curves
As the support of $D$ does not contain real points, we can write its pullback to $X_{\mathbb{C}}$ as $F+\bar{F}$ for some divisor $F$ where $\bar{F}$ denotes its complex conjugate. Let $F_{1}$ be a divisor supported on regular points fixed by the complex conjugation such that ${\mathcal{O}}_{X_{\mathbb{C}}}(F-F_{1})\in\operatorname{Pic}^{\circ}(X_{\mathbb{C}})$. Since $\operatorname{Pic}^{\circ}(X_{\mathbb{C}})$ is $2$-divisible by Lemma 2.1, there exists a divisor $G$ supported on regular points of $X_{\mathbb{C}}$ and a unit $g$ of the total quotient ring of $X_{\mathbb{C}}$ such that $2G+\operatorname{div}(g)=F-F_{1}$. This implies that
@@
Proposition 2.2 can be used to generalize a result by Geyer–Martens [geyermartens] on the structure of $2$-torsion points on a smooth real curve to singular curves. Namely, for a connected curve $X$ over $\mathbb{R}$} we consider the map
@@
Because $X$ is projective, there exists an open affine subset $U$ of $X$ which contains $X({\mathbb{R}})$ and all singular points. Indeed, if $X$ is embedded to $\mathbb{P}^{n}$, then, after applying a linear change of coordinates if necessary, we can take the principal open subset of all points where $x_{0}^{2}+\cdots+x_{n}^{2}$ is non-zero. Embedding $U$ to ${\mathbb{A}}^{m}$ gives an embedding of $X({\mathbb{R}})$ to ${\mathbb{R}}^{m}={\mathbb{A}}^{m}({\mathbb{R}})$. Let $X_{1}$ be a union of connected components of $X({\mathbb{R}})$ and $X_{2}=X({\mathbb{R}})\smallsetminus X_{1}$. Since $X_{1}$ and $X_{2}$ are disjoint compact subsets of ${\mathbb{R}}^{m}$, there is a continuous function $\psi\colon{\mathbb{R}}^{m}\to{\mathbb{R}}$ which satisfies $\psi(x)\leq-1$ for $x\in X_{1}$ and $\psi(x)\geq 1$ for $x\in X_{2}$. By the Stone–Weierstrass theorem we can uniformly approximate $\psi$ on a compact subset of ${\mathbb{R}}^{m}$ containing $X({\mathbb{R}})$ arbitrarily close by a polynomial. In particular, there exists $g\in{\mathbb{R}}[x_{1},\ldots,x_{m}]$ with $g(x)<0$ for $x\in X_{1}$ and $g(x)>0$ for $x\in X_{2}$. After replacing $g$ by a small perturbation if necessary, we can further assume that $g$ does not vanish on any singular point of $X$. We apply Proposition 2.2 to the principal divisor $D=\operatorname{div}(g)$ and obtain a divisor $E$, supported on regular points, and a unit $f$ of the total quotient ring of $X$, which is non-negative on every real point where it is defined, such that $\operatorname{div}(\frac{g}{f})=2E$. The $2$-torsion point ${\mathcal{O}}_{X}(E)$ of $\operatorname{Pic}(X)$ is then mapped by $\rm sg$ to a cohomology class represented by the function which is 1-1 on $X_{1}$ and +1+1 on $X_{2}$. Since $X_{1}$ was chosen to be an arbitrary union of connected components of $X({\mathbb{R}})$ and $X_{2}$ its complement, this shows the claim. ∎
We will see that positive $2$-torsion points are closely related to those extreme rays of $\operatorname{P}(X,L)$ with only regular zeros. Corollary 2.3 gives a way to count them.
@@ 2.4. Generalized elliptic curves
Here the first equality is a classical fact about real elliptic curves and the second equality holds by Corollary 2.3. This proves that $\dim_{{\mathbb{F}}_{2}}(\operatorname{Pic}(X)^{+}_{2})=1$. ∎
@@ 3. Proofs and examples
The dimension of the image of $H^{0}(\tilde{X}_{i},M_{i})$ under $\beta$ is bounded from above by the sum $\delta_{i}=\sum_{x\in X_{i}}\delta(x)$. Using Remark 2.7 and going through Corollary 2.6 we see that this sum is at most $1$ for irreducible $X$ and that in general $\delta_{i}\leq 2$ for every $i$. This property can also be deduced directly from the definitions by making use of the fact that $\pi^{!}_{i}\omega^{\circ}_{X}$, where $\pi_{i}\colon\tilde{X}_{i}\to X$ is the inclusion $\tilde{X}_{i}\hookrightarrow\tilde{X}$ composed with the normalization $\pi\colon\tilde{X}\to X$, is, on the one hand, the canonical sheaf of $\tilde{X}_{i}$ by [hartshorne]*Exercise III.7.2, and, on the other hand, the codifferent sheaf $\pi^{!}_{i}{\mathcal{O}}_{X}$ of $\pi_{i}$.
@@
If $X$ is smooth, then the claim follows from the Riemann–Roch theorem. Hence we assume from now on that $X$ is singular. Then each $\tilde{X}_{i}$ is isomorphic to $\mathbb{P}^{1}$ and thus the space $W$ of all $h\in H^{0}(\tilde{X}_{j},M_{j})$ with $\operatorname{div}(h)\geq\operatorname{div}(f_{j})-p-q$ has dimension $3$. By Remark 3.1 the dimension of $\beta(H^{0}(\tilde{X}_{j},M_{j}))$ is at most $2$. Since $\dim(W)=3$, there is $0\neq h\in W\cap\ker(\beta)$. Thus the claim follows if $h$ is linearly independent from $\alpha(f)$. Hence we assume in the following that $h$ and $\alpha(f)$ are linearly dependent. This can only happen if $f_{i}=0$ for all $i\neq j$ and thus $\beta(f_{j})=0$. If $r>1$, then for $g\in H^{0}(\tilde{X}_{j},M_{j})$ the condition $\beta(g)=0$ is equivalent to a vanishing condition on the points on $\tilde{X}_{j}$ that are mapped by $\pi$ to some $X_{i}$ with $i\neq j$. Since $f_{j}$ satisfies this condition and since neither $p$ nor $q$ are mapped by $\pi$ to some $X_{i}$ with $i\neq j$, the condition $\operatorname{div}(g)\geq\operatorname{div}(f_{j})-p-q$ implies that $g\in H^{0}(\tilde{X}_{j},M_{j})$ satisfies this vanishing condition as well. Thus $W$ lies in the kernel of $\beta$ which implies the claim. It remains to treat the case that $r=1$. Then by Remark 3.1 the dimension of $\beta(H^{0}(\tilde{X}_{j},M_{j}))$ is $1$ and thus the space $W\cap\ker(\beta)$ is at least of $2$-dimensional which implies the claim. ∎
Proof of Theorem 1.1.
@@
Let $X_{j}$ be an irreducible component of $X$ that contains $x$. Note that this implies $j\not\in S$. There exists $p\in\pi^{-1}(x)\cap\tilde{X}_{j}$ and we let $q=\bar{p}$ be its complex conjugate. Because $x\not\in X({\mathbb{R}})$, the points $p$ and $q$ are two distinct elements of $\tilde{X}_{j}$. Now $f$, $p$ and $q$ satisfy the assumptions of Lemma 3.2. Thus there exists $g\in H^{0}(X,L\otimes L)$ which is linearly independent of $f$ such that:
@@
Proof of Theorem 1.2.
Let $D$ be a divisor on $X$ supported on regular points such that $L\cong{\mathcal{O}}_{X}(D)$, the subsheaf of the sheaf of total quotient rings of ${\mathcal{O}}_{X}$ associated to $D$. Let $s\in H^{0}(X,L\otimes L)$ generate an extreme ray of $\operatorname{P}(X,L)$. Consider the partial normalization $\tilde{Y}\to X$ at all singular points of $X$ where $s$ vanishes. Further, let $Y\subseteq\tilde{Y}$ be the union of all irreducible components of $\tilde{Y}$ where the pullback of $s$ to $\tilde{Y}$ is not identically zero. Finally, let $\rho\colon Y\to X$ be the composition of the inclusion $Y\hookrightarrow\tilde{Y}$ with $\tilde{Y}\to X$. Then the restriction $D^{\prime}=D|_{Y}$ satisfies $\rho^{*}L\cong{\mathcal{O}}_{Y}(D^{\prime})$. We choose $f\in{\mathcal{O}}_{Y}(2D^{\prime})$, a rational function on $Y$, which corresponds to $\rho^{*}s$ under $\rho^{*}L\cong{\mathcal{O}}_{Y}(D^{\prime})$. We denote by $E_{1}$ the restriction of the zero divisor of $\rho^{*}s$ to the set of all regular points of $Y$ that lie over a singular point of $X$. Then the principal divisor of $f$ is given by $\operatorname{div}(f)=E_{1}+E_{2}-2D^{\prime}$ for some effective divisor $E_{2}$ supported on regular points. By Theorem 1.1 the divisor $E_{2}$ is a sum of real non-isolated points of $Y$. Because $f$ is non-negative on $Y({\mathbb{R}})$, it follows that $E_{2}=2F$ for some divisor $F$ because real zeros of non-negative sections must occur with even multiplicity. Consider the bilinear map
@@
The proof of Theorem 1.2 simplifies when the extreme ray $s$ has only regular zeros. In this case, the proof shows that $s$ is the image of a square under a positive semi-definite isomorphism $\psi\colon A\otimes A\to L\otimes L$ where $A$ is a line bundle. Hence we have $A=L\otimes T$ where $T\in\operatorname{Pic}(X)_{2}$ is an positive $2$-torsion point. By Lemma 2.8 there are at most two such points. However, as Example 3.5 shows, line bundles are not enough to describe extreme rays that vanish at singular points.
@@
Consider the case that $X$ is a smooth elliptic curve over $\mathbb{R}$}. A plane affine model of $X$ is given by the equation $y^{2}=x\cdot p(x)$ where $p$ is a univariate polynomial of degree $2$ with simple zeros which are either positive or non-real. Let $P$ be the point $(0,0)$ and let $O$ be the point at infinity. Then the unique (see Lemma 2.8) non-trivial positive $2$-torsion point is represented by the divisor T=POT=P-O. Indeed, we have $\operatorname{div}(x)=2T$ and $x$ is non-negative on $X({\mathbb{R}})$. Hence, if $D$ a divisor on $X$ and $L={\mathcal{O}}_{X}(D)$, then every extreme ray of $\operatorname{P}(X,L)$ is either a square of a global section of ${\mathcal{O}}_{X}(D)$ or the image of a square under the map
@@
We illustrate the proof of Theorem 1.2 in the case when $X$ is the rational curve with one cusp. A plane affine model of $X$ is given by the equation $y^{2}=x^{3}$. Let $O$ be the point at infinity and $L={\mathcal{O}}_{X}(dO)$ for $d\in{\mathbb{N}}$. Note that for $d=1,2$ the line bundle $L$ is ample but not very ample. The only possibilities for $\rho\colon Y\to X$ are the identity $X\to X$ and the normalization $\mathbb{P}^{1}\to X$. In the first case, we have $E_{1}=0$. Since the Picard group of $X$ is the additive group $\mathbb{R}$$\times{\mathbb{Z}}$, which is torsion-free, the only possibility, up to a positive scalar, for $\psi$ is the identity on $L\otimes L$. In the case $Y=\mathbb{P}^{1}$, we denote by $P$ the preimage of the cusp under $\pi$. Since $E_{2}$ has even degree and $E_{1}$ is linearly equivalent to $2dO-E_{2}$, we either have $E_{1}=2kP$ for $k=1,\ldots,d$. Here, by abuse of notation, we denote the preimage of $O$ under $\pi$ also by $O$. Again, as the Picard group of $\mathbb{P}^{1}$ is torsion-free, we find that $F$ is linearly equivalent to $dO-kP\sim(d-k)O$. In order to describe the resulting map $\psi$ on global sections explicitly, we identify the global sections of ${\mathcal{O}}_{X}(mO)$ and ${\mathcal{O}}_{Y}(mO)$ with the spaces ${\mathbb{R}}[t^{2},t^{3}]_{\leq m}$ and ${\mathbb{R}}[t]_{\leq m}$ of univariate polynomials in ${\mathbb{R}}[t^{2},t^{3}]$ and ${\mathbb{R}}[t]$ that have degree at most $m$. We get the map
@@
Theorem 1.2 gives the following nice description of the dual cone $\operatorname{P}(X,L)^{\vee}$.
@@
The proof of Theorem 1.2 shows that the description of $\operatorname{P}(X,L)^{\vee}$ in Corollary 3.6 consists of several blocks, one corresponding to each of the sheaves $A_{1},\ldots,A_{m}$ on $X$. A more careful analysis gives bounds on the size of these blocks. By the construction, the size of the block corresponding to $A_{i}$ is equal to $h^{0}(X,A_{i})$. Using the notation as in the proof of Theorem 1.2, this number is equal to $h^{0}(Y,{\mathcal{O}}_{Y}(F))$. By the Riemann–Roch theorem [stacks-project]*Lemma 0BS6 we have
@@
The proof of Corollary 1.3 is a combination of Corollary 3.6 and Proposition 2.2.
Proof of Corollary 1.3.
Because $X({\mathbb{R}})\subseteq{\mathbb{R}}^{n}=(\mathbb{P}^{n}\smallsetminus H)({\mathbb{R}})$ is compact, there exists an affine linear functional ${\mathbb{R}}^{n}\to{\mathbb{R}}$ which is positive on $X({\mathbb{R}})$. Let $\ell$ be the corresponding global section of $M={\mathcal{O}}_{X}(1)$. By slightly perturbing $\ell$ if necessary, we can assume that the zero set of $\ell$ consists only of regular points of $X$. Let $D$ be its zero divisor. By Proposition 2.2 there exists a divisor $E$, supported on regular points, and a unit $f$ of the total quotient ring of $X$, which is non-negative on every real point where it is defined, such that $D=2E+\operatorname{div}(f)$. We have isomorphisms
@@
that maps a point to the corresponding point evaluation, satisfies ${\rm ev}^{-1}(K^{\vee})=\operatorname{conv}(X({\mathbb{R}}))\subseteq{\mathbb{R}}^{n}$. Therefore, the set $\operatorname{conv}(X({\mathbb{R}}))={\rm ev}^{-1}(\psi^{\vee}(\operatorname{P}(X,L)^{\vee}))$ is a spectrahedron by Corollary 3.6. ∎
@@
It follows from Remark 3.7 that, in the situation of Corollary 1.3, the description of $\operatorname{conv}(X({\mathbb{R}}))\subseteq{\mathbb{R}}^{n}$ consists of blocks, each of size at most $\frac{n+1}{2}$. Note that in general one cannot expect a description that only involves blocks of smaller size, even when allowing a description as a spectrahedral shadow [scheiderercurves2].
@@
One checks that $X$ is a smooth elliptic curve. The hyperplane defined by $x_{0}=0$ at infinity does not intersect $X({\mathbb{R}})$. Thus Corollary 1.3 says that the convex hull of $X({\mathbb{R}})$ in the affine chart $x_{0}\neq 0$ is a spectrahedron. Indeed, it is the set of all $(x_{1},x_{2},x_{3})$ such that the matrix
@@
Therefore, it is isomorphic to the curve from Example 3.5 and ${\mathcal{O}}_{X}(1)$ is isomorphic to ${\mathcal{O}}_{X}(4O)$ where $O=(1:0:0:0)$. We use the notation as in Example 3.5. The hyperplane defined by $x_{0}=0$ at infinity does not intersect $X({\mathbb{R}})$. Thus Corollary 1.3 says that the convex hull of $X({\mathbb{R}})$ in the affine chart $x_{0}\neq 0$ is a spectrahedron. Indeed, it is the set of all $(x_{1},x_{2},x_{3})$ such that the matrix
@@
Figure 1. The (generalized) elliptic curves from Example 3.9 (left) and Example 3.10 (right).